# ZMP-HH/24-19

### v1

@@ 1 Introduction
The dark dimension scenario [1] has emerged as a result of combining Swampland principles in the context of dark energy with observational constraints, leading to a specific corner of the quantum gravity landscape [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. In particular it predicts one larger extra dimension in the micrometer range, while other possible extra dimensions are much smaller. This micron scale extra dimension is called the dark dimension. For a review, see [13] (and [14] for a different perspective). In this note we evaluate the prospects for accommodating the idea of grand unification of forces, as realized in Grand Unified Theories (GUTs), in the context of a dark dimension. We find that in this scenario, using general principles of quantum gravity and experimental observations (in particular bounds on the proton lifetime and on the mass of excited cousins of Standard Model (SM) gauge bosons) uniquely predicts a bound on the mass of the GUT gauge boson, denoted by $X$ (as a generic prediction of GUT models), of the order $M_{X}\sim 10^{15}-10^{16}\ {\rm GeV}$, and the existence of a Kaluza-Klein (KK) tower of SM gauge bosons at the mass scale of $1\sim 10\ \rm{TeV}$. Moreover we learn that the $X$ gauge boson in the 5d effective field theory (EFT) including the dark dimension is necessarily represented by a stretched solitonic string.
The aim of this paper, just as in the original work on the dark dimension scenario [1], is not to actually construct any specific models, but only to exploit general quantum gravity constraints and known experimental bounds to nevertheless arrive at concrete predictions. The reason for this point of view is twofold: First, despite constant progress (see, e.g., the reviews [15, 16] and references therein), we do not yet have all the required technology in string theory constructions to combine all the ingredients that we need at the same time and in a fully controlled way. For example we currently have no fully reliable model with complete moduli stabilization and a hierarchically small positive cosmological constant. Second, even if we found a model which satisfies all constraints, we would not know whether that is the model of our universe, because we do not know how many models exist which satisfy our constraints. Therefore any predictions based on specific models, rather than general principles, would be subject to this criticism. Instead if we only impose constraints that - to the best of our current understanding - any model would have to satisfy, we can make robust predictions. This is the path we wish to continue in this paper, by applying the lessons we have learned from quantum gravity and Swampland principles to the issue of grand unification of forces, without any specific models in mind.
@@ 3 Grand Unification and the Dark Dimension
Figure -959: Schematic representation of the dark dimension scenario. The horizontal direction represents the dark dimension of size $L_{5}$. The SM gauge fields $SU(3)\times SU(2)\times U(1)$ live in a small region of size $L_{g}$ in the dark dimension. The pink region of size $L_{m}$ denotes where the matter fields live. The string represented by the green wavy line stretched along a length of size $L_{X}$ leads to the $X$-gauge bosons.
So it seems we are at an impasse. Is there any other way to have a GUT structure at or above $10^{16}\ \rm{GeV}$? This would in particular mean that the $X$-boson mass is at $10^{16}\ {\rm GeV}$ or higher. But no localized object in 5 dimensions can have a mass higher than the Planck scale, without being a black hole! Hence the only other option is that the $X$-boson is not a point-like object, but an extended object in 5 dimensions (see Figure -959). In other words, the $X$ gauge boson must be effectively a string in the 5d sense! We are assuming that this is a point particle in 4d, as in the GUT unification picture, at least up to the GUT scale. [^2] Of course we know this cannot be strictly true in the dark dimension scenario because at an energy scale larger than $M_{5}=10^{10}\ {\rm GeV}$ we are probing the higher-dimensional Planck scale. However, we can still hope, since we want to assume the grand unification scale at $10^{16}\,{\rm GeV}$, that the 5d Planck scale, which is visible to gravitational terms in the action, is invisible to gauge theoretic interactions in the EFT description, perhaps because the gravitational Planck scale is driven by gauge neutral light states. This is for example what happens in supersymmetric gauge theories as engineered in M-theory on Calabi-Yau threefolds: Near CFT points in their moduli space, a cycle shrinks to zero size in the internal space, and even after becoming much smaller than the Planck scale, it can be unambiguously extrapolated to zero size, as clearly marked by the emergence of massless fields of an effective CFT in the strongly coupled limit. This option is of course very natural in string theory. Indeed, in brane constructions, the gauge bosons are typically obtained by stretched strings between branes. We will be assuming this is the usual realization of gauge groups as strings ending on branes. However, we will not be assuming this is a weakly coupled string (and in fact we later argue that it is strongly coupled).
@@
One may wonder whether the effective string describing the $X$-gauge boson is a weakly coupled string in stringy realizations of this scenario (see e.g. [44, 45] for orientifold realizations of GUTs and Fig. -958 for an illustration).
Figure -958: A cartoon for a would-be stringy realization where the SM gauge fields $SU(3)\times SU(2)\times U(1)$ live on the branes supported over a small region of size $L_{g}$ in the dark dimension. They intersect in a region of size $L_{m}$ where the SM matter fields are localized. The unification scale is represented by the size of the $X$-boson stretched between the branes. Such a stringy realization would be at strong coupling, so neither the branes nor their intersection region can be viewed as sharply defined, and should be taken only as a heuristic picture, which could also be hiding additional dimensions.
@@ 4 Experimental Consequences
Note that our arguments have nothing to say about the diameter of the matter locus, $L_{m}$. Indeed it is very natural that $L_{m}\ll L_{g}\sim(1-10\ \rm{TeV})^{-1}$, as the charged matter is localized in the region of intersection of the gauge branes (see Fig. -958). Thus the FCNC suppression is natural in this setup.
The thickness of the gauge boson brane will have some impact on other phenomenological aspects of the dark dimension. For example, the bound $f_{a}\lesssim{M_{5}}\sim 10^{10}\ {\rm GeV}$ for the axion decay constant in [5] was found under the assumption that the thickness of the SM brane is Planckian in 5d; here this gets modified by a factor of $\sqrt{M_{5}L_{g}}\sim 10^{3}$ and becomes $f_{a}\lesssim{M_{5}}\times 10^{3}\sim 10^{13}\ {\rm GeV}$, which sits comfortably in the allowed experimental region.
One also expects a slight modification of the cosmological scenario [2] because dark matter is produced by emission from the SM brane, which is now thicker by a factor of $M_{5}L_{g}\sim 10^{5}-10^{6}$. This leads to an enhancement in production by that factor, as can be seen by viewing the SM brane as $M_{5}L_{g}$ copies of thickness $1/M_{5}$, for which the original emissions were computed. Since the dark matter produced by emission scales with the cube of the initial temperature of the SM brane, this implies the initial temperature should be reduced by a factor of 10-100 compared to that considered in [2], to $T_{i}\sim 10-100\ \rm{MeV}$; this is still above the BBN temperature $T_{BBN}\sim\ \rm{MeV}$. It would be interesting to study this further.

### v2

@@ 1 Introduction
The dark dimension scenario [1] has emerged as a result of combining Swampland principles in the context of dark energy with observational constraints, leading to a specific corner of the quantum gravity landscape [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. In particular it predicts one larger extra dimension in the micrometer range, while other possible extra dimensions are much smaller. This micron scale extra dimension is called the dark dimension. For a review, see [21] (and [22] for a different perspective). In this note we evaluate the prospects for accommodating the idea of grand unification of forces, as realized in Grand Unified Theories (GUTs), in the context of a dark dimension. We find that in this scenario, using general principles of quantum gravity and experimental observations (in particular bounds on the proton lifetime and on the mass of excited cousins of Standard Model (SM) gauge bosons) uniquely predicts a bound on the mass of the GUT gauge boson, denoted by $X$ (as a generic prediction of GUT models), of the order $M_{X}\sim 10^{15}-10^{16}\ {\rm GeV}$, and the existence of a Kaluza-Klein (KK) tower of SM gauge bosons at the mass scale of $1\sim 10\ \rm{TeV}$. Moreover we learn that the $X$ gauge boson in the 5d effective field theory (EFT) including the dark dimension is necessarily represented by a stretched solitonic string.
The aim of this paper, just as in the original work on the dark dimension scenario [1], is not to actually construct any specific models, but only to exploit general quantum gravity constraints and known experimental bounds to nevertheless arrive at concrete predictions. The reason for this point of view is twofold: First, despite continuous progress (see, e.g., the reviews [23, 24] and references therein), we do not yet have all the required technology in string theory constructions to combine all the ingredients that we need at the same time and in a fully controlled way. For example we currently have no fully reliable model with complete moduli stabilization and a hierarchically small positive cosmological constant. Second, even if we found a model which satisfies all constraints, we would not know whether that is the model of our universe, because we do not know how many models exist which satisfy our constraints. Therefore any predictions based on specific models, rather than general principles, would be subject to this criticism. Instead if we only impose constraints that - to the best of our current understanding - any model would have to satisfy, we can make robust predictions. This is the path we wish to continue in this paper, by applying the lessons we have learned from quantum gravity and Swampland principles to the issue of grand unification of forces, without any specific models in mind.
@@ 3 Grand Unification and the Dark Dimension
Figure -1019: Schematic representation of the dark dimension scenario. The horizontal direction represents the dark dimension of size $L_{5}$. The SM gauge fields $SU(3)\times SU(2)\times U(1)$ live in a small region of size $L_{g}$ in the dark dimension. The pink region of size $L_{m}$ denotes where the matter fields live. The string represented by the green wavy line stretched along a length of size $L_{X}$ leads to the $X$-gauge bosons.
So it seems we are at an impasse. Is there any other way to have a GUT structure at or above $10^{16}\ \rm{GeV}$? This would in particular mean that the $X$-boson mass is at $10^{16}\ {\rm GeV}$ or higher. But no localized object in 5 dimensions can have a mass higher than the Planck scale, without being a black hole! Hence the only other option is that the $X$-boson is not a point-like object, but an extended object in 5 dimensions (see Figure -1019). In other words, the $X$ gauge boson must be effectively a string in the 5d sense! We are assuming that this is a point particle in 4d, as in the GUT unification picture, at least up to the GUT scale. [^2] Of course we know this cannot be strictly true in the dark dimension scenario because at an energy scale larger than $M_{5}=10^{10}\ {\rm GeV}$ we are probing the higher-dimensional Planck scale. However, we can still hope, since we want to assume the grand unification scale at $10^{16}\,{\rm GeV}$, that the 5d Planck scale, which is visible to gravitational terms in the action, is invisible to gauge theoretic interactions in the EFT description, perhaps because the gravitational Planck scale is driven by gauge neutral light states. This is for example what happens in supersymmetric gauge theories as engineered in M-theory on Calabi-Yau threefolds: Near CFT points in their moduli space, a cycle shrinks to zero size in the internal space, and even after becoming much smaller than the Planck scale, it can be unambiguously extrapolated to zero size, as clearly marked by the emergence of massless fields of an effective CFT in the strongly coupled limit. This is also compatible with the fact that for the EFT in 4d $M_{GUT}<M_{pl}$ and the 5d Planck scale is hard to detect from the 4d perspective [48].
This option is of course very natural in string theory. Indeed, in brane constructions, the gauge bosons are typically obtained by stretched strings between branes. We will be assuming this is the usual realization of gauge groups as strings ending on branes. However, we will not be assuming this is a weakly coupled string (and in fact we later argue that it is strongly coupled).
@@
One may wonder whether the effective string describing the $X$-gauge boson is a weakly coupled string in stringy realizations of this scenario (see e.g. [53, 54] for orientifold realizations of GUTs and Fig. -1018 for an illustration).
Figure -1018: A cartoon for a would-be stringy realization where the SM gauge fields $SU(3)\times SU(2)\times U(1)$ live on the branes supported over a small region of size $L_{g}$ in the dark dimension. They intersect in a region of size $L_{m}$ where the SM matter fields are localized. The unification scale is represented by the size of the $X$-boson stretched between the branes. Such a stringy realization would be at strong coupling, so neither the branes nor their intersection region can be viewed as sharply defined, and should be taken only as a heuristic picture, which could also be hiding additional dimensions.
@@ 4 Experimental Consequences
Note that our arguments have nothing to say about the diameter of the matter locus, $L_{m}$. Indeed it is very natural that $L_{m}\ll L_{g}\sim(1-10\ \rm{TeV})^{-1}$, as the charged matter is localized in the region of intersection of the gauge branes (see Fig. -1018). Thus the FCNC suppression is natural in this setup.
The thickness of the gauge boson brane will have some impact on other phenomenological aspects of the dark dimension. For example, the bound $f_{a}\lesssim{M_{5}}\sim 10^{10}\ {\rm GeV}$ for the axion decay constant in [5] was found under the assumption that the thickness of the SM brane is Planckian in 5d; here this gets modified by a factor of $\sqrt{M_{5}L_{g}}\sim 10^{3}$ and becomes $f_{a}\lesssim{M_{5}}\times 10^{3}\sim 10^{13}\ {\rm GeV}$, which sits comfortably in the allowed experimental region. It would also be interesting to study the impact of this scenario on the cosmology of the dark dimension scenario [2].